OpenStudy (anonymous):

Use mathematical induction to prove that the statement is true for every positive integer n. 2 is a factor of n^2 - n +2

OpenStudy (anonymous):

P(n)=n^2-n+2 P(1)=1^2-1+2=0,which is a multiple of 2 P(1) is true. Assume that P(k) is true. P(k)=k2-k+2=2t (say) k^2=2t+k-2 P(k+1)=(k+1)^2-(k+1)+2 =k^2+2k+1-k-1+2 =K^2+k+2 =2t+k-2+k+2 =2t+2k =2(t+k) =2*an integer Hence P(k+1) is true. By induction P(n) is true for all n (positive integer).

OpenStudy (anonymous):

Thank you yeah i just figured some of this out i have another one i need help on